Volume/Time - IV mL Rate Questions

A.Given a certain amount of liquid and a time period, what is the necessary IV flow rate in mL/hr? Measurement used when IV regulated electronically by infusion pump.

Formula:

Volume (mL)

Time (hr)

 =  Y (Flow Rate in mL/hr)

Example: Infuse 250 mL over the next 120 minutes by infusion pump.

Volume (mL)

Time (hr)

 =  Y (Flow Rate in mL/hr)

B.Convert 120 minutes to hours.

min → hr    ( ÷ by 60 )

120 min ÷ 60 = 2 hr

250 mL

2 hr

 =  125 mL/hr

Example: Ordered 1000 mL D5W IV to infuse in 10 hours by infusion pump.

Volume (mL)

Time (hr)

 =  Y (Flow Rate in mL/hr)

1000 mL

10 hr

 =  100 mL/hr

Example: Convert 5,000 mcg to mg.

mcg → mg → g → kg    ( ÷ by 1,000 )

5,000 mcg ÷ 1,000 = 5 mg

Example: Convert 44 lb to kg.

lb → kg    ( ÷ by 2.2 )

44 lb ÷ 2.2 = 20 kg

Example: Convert 0.003 L to mcL.

mcL ← mL ← L ← kL    ( x by 1,000 )

0.003 L x 1,000 = 3 mL

3 mL x 1,000 = 3,000 mcL

Example: Convert 5 hours to minutes.

min ← hr    ( x by 60 )

5 hr x 60 = 300 min

1.Question: Ordered 270 mL NS IV to infuse in 5.5 hr by infusion pump. What is the IV flow rate in mL/hr?

2.Question: Potassium chloride is available as 0.013 kg per tablet. Potassium Chloride (K-Dur), 26 mg, is ordered. How many tablets would the nurse administer?

3.Give patient 24.9 mg of dopamine in 283 mL of D5W to be infused at a rate of 17 mg/hr. Calculate the flow rate in mL/hr.

4.Give patient 24.9 mg of dopamine in 283 mL of D5W to be infused at a rate of 17 mg/hr. Calculate the flow rate in mL/hr.
Answer #1

1.Question: Ordered 270 mL NS IV to infuse in 5.5 hr by infusion pump. What is the IV flow rate in mL/hr?

Rate of IV fluid = volume/ time

                           = 270/5.5 = 49ml/hr

2.Question: Potassium chloride is available as 0.013 kg per tablet. Potassium Chloride (K-Dur), 26 mg, is ordered. How many tablets would the nurse administer?

One tablet of Potassium Chloride = 0.013 Kg =13 mg

The dose to be given is 26 mg

26 mg contains how many tablets ?

That is 26/13 = 2 tablets

3.Give patient 24.9 mg of dopamine in 283 mL of D5W to be infused at a rate of 17 mg/hr. Calculate the flow rate in mL/hr.

Rate of flow = volume / time

                   

24.9 mg in 283ml of D5W, then 17mg is in how much ml?

24.9 283 17 x

x is the volume o fluid containing 17 mg Dopamine.

Therefore,

     x =17×283/24.9 = 193.2 ml

193 ml / hr is the flow rate to be set.

4.Give patient 24.9 mg of dopamine in 283 mL of D5W to be infused at a rate of 17 mg/hr. Calculate the flow rate in mL/hr.

Please check the answer for question number 3